OpenStudy (anonymous):

a cash box contained 14.55 in quarters,D,N. if there were 3 more than twice as many D as N and three less than three times as many Q as N how many of each does she have?

OpenStudy (anonymous):

help

OpenStudy (radar):

D=number of dimes N=number of nickels Q=number of quarters N=3+2D N=3Q-3 5N+25Q+10D=1455 Have I got this right? !!!!

OpenStudy (radar):

If so there is three equations with 3 unknowns and can be solved.

OpenStudy (radar):

Somewhere I erred in the translation. If there were 3 more than twice as many dimes as nickels. What is that saying? N=2D + 3 ??? Someone help ISA DIA

OpenStudy (radar):

With the relationships I had in my first post, I come out with a fraction for the number of nickels.....I know that can't be correct. Are you there robtobey??

OpenStudy (anonymous):

A Mathematica solution is attached.

OpenStudy (radar):

Thanks robtobey, hopefully ISA DIA is still interested.